find the vertex and axis of symmetry of the graph of the function
f(x)= xsquared + 4x + 4

Answer by Alan3354(69443)

f(x)= x^2 + 4x + 4
Vertex at -b/2a = -4/2 = -2 (x value)
f(-2) = 0
Vertex at (-2,0)
-----------------
Axis of symmetry is x = -2